OpenStudy (anonymous):

I need help solving these 3 equations please! Simplify. 1. |-4b-8|+|-1-b^2|+2b^3 , b=-2 2. -(4h-1)/1+h , h=2 3. 9c-6d+4c-3d

OpenStudy (owlcoffee):

Well, let's begin with the first one: \[\left| -4b-8 \right|+\left| -1-b^2\right|+2b^3 \] And since we are given the value of the variable, we can simply plug it in: \[\left| -4(-2)-8 \right|+\left| -1-(-2)^2\right|+2(-2)^3 \]

OpenStudy (anonymous):

Where would I start? This problem is so confusing, i know how to do absolute value inequalities but nothing this complicated.

OpenStudy (owlcoffee):

With the correct operations, you'll obtain: \[\left| 8-8 \right|+\left| -1-4 \right|-16\] So, all you have to do is operate further.

OpenStudy (anonymous):

I feel really stupid right now trying to figure that out please help. :( So how would I operate further?

OpenStudy (owlcoffee):

Sure: \[\left| 8-8 \right|+\left| -1-4 \right|-16 = \left| 0 \right|+\left| -5 \right|-16\] \[\left| 0 \right|+\left| -5 \right|-16=0+5+16\]